OpenStudy (anonymous):

change 9x squared + 4y squared + 126x + 64y + 661=0 into standared form

OpenStudy (anonymous):

I am trying to learn to do them of course, but once I do them am not sure if I have the right answer....just getting the answers would not help me if I can't do them myself

OpenStudy (anonymous):

Not to worry

OpenStudy (anonymous):

I will surely help you

OpenStudy (anonymous):

Thank you so much balaa

OpenStudy (anonymous):

Here is one to get you started 9x^2+4y^2+126x+64y+661=0 9x^2 + 126x + 4y^2 + 64y = -661 9(x^2 + 14x) + 4(y^2 + 16y) = -661 9(x + 7)^2 + 4(y + 8)^2 = -661 + 9•7^2 + 4•8^2 9(x + 7)^2 + 4(y + 8)^2 = 36 9(x + 7)^2 / 36 + 4(y + 4)^2 / 36 = 36/36 (x + 7)^2 / 4 + (y + 4)^2 / 9 = 1

OpenStudy (anonymous):

Standard form (x + 7)^2 / 4 + (y + 4)^2 / 9 = 1

OpenStudy (anonymous):

thank you blaa I think I'm getting it